Status Report

NASA Space Station On-Orbit Status 9 June 2003

By SpaceRef Editor
June 9, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 9 June 2003

All ISS systems continue to function nominally, except as noted previously or below.  Week 6 of Increment 7 is underway.

In preparation for Progress 11P docking on Wednesday morning (6/11), CDR Yuri Malenchenko and FE/SO Ed Lu performed a 2-hr. refresher training on the TORU backup docking control system, involving the review of pertinent on-board data files and the uplinked results of rendezvous math modeling, as well as a tagup with a TORU instructor.  [The teleoperated TORU system allows manual control of Progress cargo ship docking.  If the primary automatic KURS system should fail, Malenchenko would use the TV picture of the approaching ISS as seen from the Klest-M video camera aboard Progress, and a hand controller that provides steering signals to the Progress by radio link.  The on-board training syllabus for all rendezvous approach and docking phases includes off-nominal situations with failed comm channels and loss of video picture.]

Working from the Russian ISS Wiener laptop, Malenchenko performed the second stage of the loaded a software upgrade (to vers. 1.2) for the multiplex bus synchronization unit of the Russian OpsLAN network (SUBA BSMM) from a CD-ROM.  Stage 1 of the upgrade was completed by Nikolai Budarin on 4/24 (after his first attempt, at 4/9, had failed).

Also in preparation for the Progress docking, the crew configured the Russian segment (RS) Soyuz LIV video system to the US segment (USOS) Ku-band assets for live TV coverage of the docking.  The transmission setup was then checked out in a stand-alone test.  Ed Lu concluded the activities by taking a video survey of the TV “drag-thrus”, i.e., the cables routed temporarily through the hatches, for ground inspection.

After having switched filter bed 1 of the SM’s harmful impurities removal unit (BMP) to Regeneration cycle yesterday, Malenchenko today terminated the bake-out, moding the channel back to Purify.  Later, he initiated regeneration on filter channel 2.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]

Yuri conducted his regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) payload which experiments with growth and development of plants under spaceflight conditions in the Lada-2 greenhouse.

Later, Malenchenko dismantled the Molniya-SM/LSO payload at SM window #3, deactivating the French-provided EGE1 laptop.  He then stowed all hardware.  [Objective of Molniya-SM, which Yuri installed yesterday as a task list item, was to record storm phenomena and other related events in the Earth’s equatorial regions, similar to the French LSO experiment, which looked for glow phenomena, called “sprites”, sometimes observable above thunderstorm clouds.]

In support of the Russian Diatomeya photo and video imagery of bioproductive areas of the world’s oceans, Malenchenko today focused on the Caribbean and Sargasso seas.  [Yuri used the Nikon-F5 digital still camera with f80mm focal length lens and a DVCAM 150 digital video system, storing on a cassette.  Important targets, from the oceanography standpoint, were sea bloom features (color-contrasting features, TsKO) on the surface of the ocean (colored spots, bands, swirls), and pronounced irregularities of the cloud cover along the track.]

At 2:00pm EDT, the Vozdukh carbon dioxide (CO2) removal system was switched from manual mode 3/5 to automatic mode.  [In automatic mode, the Vozdukh is controlled by the SM’s BVS onboard computer complex in conjunction with the BA automatic control unit, operating in closed-loop control with the IK0501 gas analyzer.  The IK0501 was successfully recalibrated by MCC-Moscow on 6/2 in a 45 min. activity which brought its CO2 partial pressure readings into close match with those of the U.S. MCA (major constituent analyzer).]

The Science Officer completed the daily routine maintenance of the SM’s SOZh environment control & life support system as well as the regular status checkup of the autonomous :Lab payloads, while the CDR prepared the daily IMS inventory “delta” file.

Both crewmembers conducted their regular daily physical exercise program of 2.5 hrs., and Ed Lu completed the weekly and monthly maintenance of the TVIS treadmill.  [Replacement straps for the degraded TVIS crew harness could not be loaded on the Progress 11P before its launch on Sunday, 6/8.  The ground will work with the crew to troubleshoot the equipment to extend its service life as much as possible.  First task for the crew, in the “job jar”: provide the ground with photo and video documentation of the damaged harness hardware.]

Moscow reported that SAM battery #4 was successfully put back on line and is now working nominally at 80 amp-hours.  This restores the full complement of eight 800A batteries in the SM (but according to TsUP, “it remains to be seen how #4 behaves”).

Progress 11P continues its slow catch-up approach to the ISS, with all systems reported to be nominal.  A third (and final) maneuver burn (DV3, with 3 m/s delta-V) is scheduled for tomorrow morning at 7:45am EDT.  Docking at the DC-1 airlock compartment is set for 7:17am EDT on 6/11.

Saturday’s successful camera test for the Shuttle Viewing DTO (development test objective) was judged highly satisfactory at MCC-H.  [Results of the tests are presently being evaluated.  They used the MISSE (Materials ISS Experiment) at the Airlock endcone and the starboard EVA toolbox #1 as targets simulating observing/mapping conditions of a Shuttle orbiter’s TPS tiles and RCC (reinforced carbon-carbon) elements on orbit.]

Today’s optional CEO targets, no longer limited in the current LVLH attitude and including the targets of the Lewis & Clark 200-year memorial locations, were Hyderabad, India (this Indian megacity lies in the Deccan Plateau, inland about 200 miles west northwest of the delta of the Godavari River.  As the crew approached from the SW, they were advised to look right of track), Recife, Brazil (the ISS pass was probably early enough in the day to limit fair weather cumulus development.  Crew had a near-nadir view of South America’s easternmost city on the coast), Rome, Italy (the “Eternal City” lies just inland, and was slightly to the right of track), Lake Poopo (as the El Niño rapidly wanes, CEO will continue to monitor conditions in this lake and adjacent salars), Los Angeles, California (the crew had a very nice pass over the LA basin, just left of track), Las Vegas, Nevada (Vegas still will be sizzling at this hour, but with clear weather and a nadir view), and the Mouth of Bad River, SD (LEWIS & CLARK SITE: This small river joins the Missouri River from the west, just opposite the city of Pierre and about 7 miles south of the dam for large Lake Oahe.  Late afternoon thunderstorms were probably east of this area by the time of the ISS pass).
CEO images can be viewed at the websites
See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:15pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (20 amp mode).  Vozdukh CO2 scrubber is On (automatic mode).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Regeneration mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 21.7.
  • Node:  Pressure (mmHg) — 743.25; temperature (deg C) — 22.4 (shell); ppO2 (mmHg) — 171.4; ppCO2 (mmHg) — 4.2.
  • U.S. Lab:  Pressure (mmHg) — 745.22; temperature (deg C) — 23.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):  Pressure (mmHg) — 745.32; temperature (deg C) — 25.3; shell heater temp (deg C) — 22.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) –n/a.
  • PMA-1:  Shell heater temp (deg C) — 23.5.
  • PMA-2:  Shell heater temp (deg C) — 20.3.

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available: 3478 kg (7668 lb) as of 6/5 [SM(774) + FGB(2547) + Progress(182)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in fixed dual angle mode, with “night glider” drag reduction.
  • SM batteries:  All batteries (8) are in “Partial Charge” mode.
  • FGB batteries:  Battery #4 is disconnected; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational. Being upgraded.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational. Being upgraded.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: -10 deg, pitch: -7.2 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.
  • Solar Beta angle:  0.8 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally [IAC-1 (internal audio controller #1) being analyzed after self-test error.  IAC-2 is prime.]
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched at WS4, with KA power. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:10am EDT [= epoch]):

  • Mean altitude — 389.5 km
  • Apogee — 393.9 km
  • Perigee — 385.1 km
  • Period — 92.35 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.000655
  • Orbits per 24-hr. day — 15.59
  • Solar Beta Angle — 0.8 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 25980
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.